# On the Statistical Complexity of Estimating VENDI Scores from Empirical Data

### v1

@@ 3.2 Matrix-based Entropy Functions and VENDI Score
Given a PSD matrix $A\in\mathbb{R}^{d\times d}$ with a unit trace $\mathrm{Tr}(A)=1$, $A$’s eigenvalues form a probability model. The order-$\alpha$ Renyi entropy of matrix $A$ is defined using the order-$\alpha$ entropy of its eigenvalues as
@@
Specifically, in the case of $\alpha=1$, the above definition reduces to the Shannon entropy of the eigenvalues as $H_{1}(A)\,:=\,\sum_{i=1}^{d}\lambda_{i}\log({1}/{\lambda_{i}})$. Applying the above standard definitions to the normalized kernel matrix $\frac{1}{n}K$, [10] define the order-$\alpha$ VENDI score for samples $x_{1},\ldots,x_{n}$ as
@@
We note that in the case of $\alpha=2$, the definition of $\mathrm{VENDI}_{2}$ is identical to the RKE score proposed by [2]. In this particular case, the score can be formulated using the Frobenius norm of the kernel matrix, which we denote by $\|\cdot\|_{F}$,
$\mathrm{RKE}(x_{1},\ldots,x_{n})\,=\,\mathrm{VENDI}_{2}(x_{1},\ldots,x_{n})\,=\,\Bigl{\|}\frac{1}{n}K\Bigr{\|}^{-2}_{F}$
3.3 Statistical Interpretation of VENDI Score
To derive the VENDI statistic that is supposed to be estimated by the VENDI score, we review the following discussion from [23, 1, 2]. First, note that the normalized kernel matrix $\frac{1}{n}K$, whose eigenvalues are used in the definition of VENDI score, can be written as:
@@ 3.3 Statistical Interpretation of VENDI Score
where $\Phi\in\mathbb{R}^{n\times d}$ is an $n\times d$ matrix whose rows are the feature presentations of samples, i.e., $\phi(x_{1}),\ldots,\phi(x_{n})$. Therefore, the normalized kernel matrix $\frac{1}{n}K$ shares the same non-zero eigenvalues with $\frac{1}{n}\Phi^{\top}\Phi$, in which the multiplication order is flipped. Note that $\frac{1}{n}\Phi^{\top}\Phi$ is defined as the empirical kernel covariance matrix $C_{X}$:
$C_{X}:=\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})\phi(x_{i})^{\top}=\frac{1}{n}\Phi^{\top}\Phi.$
As a result, the empirical covariance matrix $C_{X}$ and kernel matrix $\frac{1}{n}K$ have the same non-zero eigenvalues and therefore share the same matrix-based entropy value for any order $\alpha$: $H_{\alpha}(\frac{1}{n}K)=H_{\alpha}(C_{X})$. Therefore, if we consider the underlying kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P_{X}}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$, we can define the VENDI statistic as follows.
@@
Given data distribution $P_{X}$, we define the order-$\alpha$ VENDI statistic, ${\mathrm{VENDI}}_{\alpha}(P_{X})$, using the matrix-based entropy of the underlying kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P_{X}}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$ as
@@
In the next sections, we study the complexity of estimating the above VENDI statistic from a limited number of samples.
4 Statistical Convergence of Vendi Scores in Finite-Dimension Kernels
Given the definitions of the VENDI score and the VENDI statistic, a relevant question is how many samples are required to accurately estimate the VENDI statistic using the VENDI score. To address this question, we first prove the following concentration bound on the vector of ordered eigenvalues $[\lambda_{1},\ldots,\lambda_{n}]$ of the kernel matrix for a normalized kernel function. We defer the proof of the theoretical results to the Appendix.
@@ 4 Statistical Convergence of Vendi Scores in Finite-Dimension Kernels
Consider a normalized kernel function $k$ satisfying $k(x,x)=1$ for every $x\in\mathcal{X}$. Let $\widehat{\boldsymbol{\lambda}}_{n}$ be the vector of sorted eigenvalues of the normalized kernel matrix $\frac{1}{n}K$ for $n$ independent samples $x_{1},\ldots,x_{n}\sim P_{X}$. If we define $\widetilde{\boldsymbol{\lambda}}$ as the vector of sorted eigenvalues of underlying covariance matrix $\widetilde{C}_{X}$, then for every $\delta\geq\exp\bigl{(}(2-n)/8\bigr{)}$, the following inequality holds with probability at least $1-\delta$:
@@
The above theorem implies the following corollary on a dimension-free convergence guarantee for order-$\alpha$ VENDI score with $\alpha\geq 2$.
@@
In the setting of Theorem 1, for every $\alpha\geq 2$ and $\delta\geq\exp\bigl{(}(2-n)/8\bigr{)}$, the following bound holds with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{VENDI}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}\leq\sqrt{\frac{32\log\bigl{(}2/\delta\bigr{)}}{n}}$
On the other hand, if the order $\alpha$ of VENDI score is less than $2$, then Theorem 1 provides a statistical convergence guarantee under a bounded dimension kernel map as shown in the following corollary.
@@
In the setting of Theorem 1, consider a finite dimension kernel map where $\mathrm{dim}(\phi)=d<\infty$. Then, for every $21\leq\alpha<$ 2 and $\delta\geq\exp\bigl{(}(2-n)/8\bigr{)}$, the following bound holds with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{VENDI}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}\leq\sqrt{\frac{32d^{2-\alpha}\log\bigl{(}2/\delta\bigr{)}}{n}}$
Therefore, assuming a bounded feature map $d<\infty$, the above result indicates the convergence of the VENDI score to the underlying VENDI statistic given $n=O(d)$ samples. We note that the result is consistent with our numerical observations of the convergence of VENDI score using a finite dimension kernel, e.g. the cosine similarity kernel (Figure 1). Next, we discuss how to extend the above result to an infinite-dimension kernel map by defining the truncated VENDI statistic.
5 Truncated Vendi Statistic and its Estimation via Proxy Kernels
Corollaries 1,2 demonstrate that if the VENDI score order $\alpha$ is greater than $2$ or the kernel feature map dimension $d$ is finite, then the VENDI score can converge to the underlying VENDI statistic with $n=O(d)$ samples. However, the theoretical results do not apply to an order $21\leq\alpha<$ 2 when the kernel map dimension is infinite, e.g. the original order-1 VENDI score [1] with a Gaussian kernel. Our numerical observations indicate that a standard sample size below 20000 could be insufficient for the convergence of order-1 VENDI score (Figure 1). To address this gap, here we define the truncated VENDI statistic and then show the existing kernel approximation algorithms for VENDI score concentrate around this modified statistic.
@@ 5 Truncated Vendi Statistic and its Estimation via Proxy Kernels
Consider data distribution $P$ and its underlying kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$. Then, for parameter $t\geq 1$, consider the top-$t$ eigenvalues of $\widetilde{C}_{X}$: $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{t}$. Define $S_{t}=\sum_{i=1}^{t}\lambda_{i}$, and consider the probability sequence $\widetilde{\lambda}_{i}=\lambda_{i}+\frac{1-S_{t}}{t}$ for $i=1,\ldots,t$. Then, we define the order-$\alpha$ $t$-truncated VENDI statistic as
$\mathrm{VENDI}_{\alpha}^{(t)}(P_{X}):=\exp\Bigl{(}\frac{1}{1-\alpha}\log\Bigl{(}\sum_{i=1}^{t}{\widetilde{\lambda}}^{\alpha}_{i}\Bigr{)}\Bigr{)}.$
According to the above definition, we find the probability model with the minimum $\ell_{2}$-norm difference from the $t$-dimensional vector $[\lambda_{1},\ldots,\lambda_{t}]$ including only the top-$t$ eigenvalues. Then, we use the order-$\alpha$ entropy of the probability model to define the order-$\alpha$ $t$-truncated VENDI statistic. Our next result shows that this statistic can be generally estimated using $n=O(t)$ samples.
@@
Consider the setting in Theorem 1. Then, for every $\delta\geq\exp((2-n)/8)$, the difference between the $t$-truncated VENDI statistic and its empirical estimation from samples $x_{1},\ldots,x_{n}$ is bounded as follows with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}\leq\sqrt{\frac{32\max\{1,t^{2-\alpha}\}\log\bigl{(}2/\delta\bigr{)}}{n}}$
As implied by Theorem 2, the $t$-truncated VENDI statistic can be estimated using $O(t)$ samples, i.e. the truncation parameter $t$ plays the role of the bounded dimension of a finite-dimension kernel map. Our next theorem shows that the Nyström method used by [10] and the FKEA method proposed by [4] to reduce the computational complexity of the VENDI score are expected to have a bounded distance to the truncated VENDI statistic.
@@
Consider the setting of Theorem 1. (a) Assume that the kernel function is shift-invariant and the FKEA method with $t$ random Fourier features is used to approximate the VENDI score. Then, for every $\delta\geq\exp((2-n)/8)$, with probability at least $1-\delta$:
$\Bigl{|}\mathrm{FKEA}\text{-}\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\sqrt{\frac{128\max\{1,t^{2-\alpha}\}\log\bigl{(}3/\delta\bigr{)}}{\min\{n,t\}}}$
@@
$\Bigl{|}\mathrm{Nystrom}\text{-}\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\Bigr{|}\leq\mathcal{O}\Bigl{(}\sqrt{\frac{\max\{1,t^{2-\alpha}\}\log\bigl{(}2/\delta\bigr{)}t\tau^{2}\log(n)^{2}}{n}}\Bigr{)}$
Theorem 3 shows that the application of the fast kernel approximation algorithms in FKEA and Nyström with parameter $t$ will provide an estimation of the $t$-truncated VENDI statistic, connecting the defined statistic to the existing algorithms for computing the VENDI score.

### v3

@@ 3.2 Matrix-based Entropy Functions and Vendi Score
For a PSD matrix $A\in\mathbb{R}^{d\times d}$ with unit trace $\mathrm{Tr}(A)=1$, $A$’s eigenvalues form a probability model. The order-$\alpha$ Renyi entropy of matrix $A$ is defined using the order-$\alpha$ entropy of its eigenvalues as
@@
For the special case $\alpha=2$, one can consider the Frobenius norm $\|\cdot\|_{F}$ and apply the identity $\bigl{\|}A\bigr{\|}_{F}^{2}=\sum_{i=1}^{d}\lambda^{2}_{i}$ to show $H_{2}(A)=\log\bigl{(}1/\bigl{\|}A\bigr{\|}_{F}^{2}\bigr{)}$. Moreover, for $\alpha=1$, the above definition reduces to the Shannon entropy of the eigenvalues as $H_{1}(A)\,:=\,\sum_{i=1}^{d}\lambda_{i}\log({1}/{\lambda_{i}})$ [Rényi, 1961].
Jalali et al. [2023] applies the above definition for order $\alpha=2$ to the normalized kernel similarity matrix $\frac{1}{n}K$ to define the RKE diversity score (called RKE mode count), which reduces to
$\mathrm{RKE}(x_{1},\ldots,x_{n})\,:=\,\exp\Bigl{(}H_{2}\bigl{(}\frac{1}{n}K\bigr{)}\Bigr{)}=\,\Bigl{\|}\frac{1}{n}K\Bigr{\|}^{-2}_{F}$ (7)
For a general entropy order $\alpha$, [Friedman and Dieng, 2023, Pasarkar and Dieng, 2024] apply the matrix-based entropy definition to the normalized kernel matrix $\frac{1}{n}K$ and define the order-$\alpha$ Vendi score for samples $x_{1},\ldots,x_{n}$ as
@@
Specifically, for order $\alpha=1$, the above definition results in the standard (order-1) Vendi score in Equation (2).
3.3 Statistical Analysis of Vendi Score
To derive the population limits of Vendi and RKE scores under infinite sampling, which we call population Vendi and population RKE, respectively, we review the following discussion from [Bach, 2022, Jalali et al., 2023]. First, note that the normalized kernel matrix $\frac{1}{n}K$, whose eigenvalues are used in the definition of Vendi score, can be written as:
@@ 3.3 Statistical Analysis of Vendi Score
where $\Phi\in\mathbb{R}^{n\times d}$ is an $n\times d$ matrix whose rows are the feature presentations of samples, i.e., $\phi(x_{1}),\ldots,\phi(x_{n})$. Therefore, the normalized kernel matrix $\frac{1}{n}K$ shares the same non-zero eigenvalues with $\frac{1}{n}\Phi^{\top}\Phi$, where the multiplication order is flipped. Note that $\frac{1}{n}\Phi^{\top}\Phi$ is equal to the empirical kernel covariance matrix $\widehat{C}_{X}$ defined as:
$\widehat{C}_{X}:=\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})\phi(x_{i})^{\top}=\frac{1}{n}\Phi^{\top}\Phi.$
As a result, the empirical covariance matrix $\widehat{C}_{X}=\frac{1}{n}\Phi^{\top}\Phi$ and kernel matrix $\frac{1}{n}K=\frac{1}{n}\Phi\Phi^{\top}$ share the same non-zero eigenvalues and therefore have the same matrix-based entropy value for every order $\alpha$: $H_{\alpha}(\frac{1}{n}K)=H_{\alpha}(\widehat{C}_{X})$. Therefore, if we consider the population kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P_{X}}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$, we can define the population Vendi score as follows.
@@
Given data distribution $P_{X}$, we define the order-$\alpha$ population Vendi, ${\mathrm{Vendi}}_{\alpha}(P_{X})$, using the matrix-based entropy of the population kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P_{X}}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$ as
@@
Note that the population RKE score is identical to the population $\mathrm{Vendi}_{2}$, since RKE and $\mathrm{Vendi}_{2}$ are the same.
4 Statistical Convergence of Vendi and RKE Scores
Given the definitions of the Vendi score and the population Vendi, a relevant question is how many samples are required to accurately estimate the population Vendi using the Vendi score. To address this question, we first prove the following concentration bound on the vector of ordered eigenvalues $[\lambda_{1},\ldots,\lambda_{n}]$ of the kernel matrix for a normalized kernel function.
@@ 4 Statistical Convergence of Vendi and RKE Scores
Consider a normalized kernel function $k$ satisfying $k(x,x)=1$ for every $x\in\mathcal{X}$. Let $\widehat{\boldsymbol{\lambda}}_{n}$ be the vector of sorted eigenvalues of the normalized kernel matrix $\frac{1}{n}K$ for $n$ independent samples $x_{1},\ldots,x_{n}\sim P_{X}$. If we define $\widetilde{\boldsymbol{\lambda}}$ as the vector of sorted eigenvalues of underlying covariance matrix $\widetilde{C}_{X}$, then if $n\geq 2+8\log(1/\delta)$, the following inequality holds with probability at least $1-\delta$:
@@
Proof. We defer the proof to the Appendix.
Theorem 1 results in the following corollary on a dimension-free convergence guarantee for every $\mathrm{Vendi}_{\alpha}$ score with order $\alpha\geq 2$, including the RKE score (i.e. $\mathrm{Vendi}_{2}$).
@@
In the setting of Theorem 1, for every $\alpha\geq 2$ and $n\geq 2+8\log(1/\delta)$, the following bound holds with probability at least $1-\delta$:
$\Bigl{|}\mathrm{Vendi}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{Vendi}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\Bigr{|}\leq$ $\sqrt{\frac{32\log\frac{2}{\delta}}{n}}$
Notably, for $\alpha=2$, we arrive at the following bound on the gap between the empirical and population RKE scores:
$\Bigl{|}\mathrm{RKE}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{-1/2}-\mathrm{RKE}\bigl{(}P_{X}\bigr{)}^{-1/2}\Bigr{|}\leq\sqrt{\frac{32\log\frac{2}{\delta}}{n}}$
Proof. We defer the proof to the Appendix. Therefore, the bound in Corollary 1 holds regardless of the dimension of kernel feature map, indicating that the RKE score enjoys a universal convergence guarantee across all kernel functions. Next, we show that Theorem 1 implies the following corollary on a dimension-dependent convergence guarantee for order-$\alpha$ Vendi score with $21\leq\alpha<$ 2, including standard (order-$1$) Vendi score.
@@
In the setting of Theorem 1, consider a finite dimension kernel map where we suppose $\mathrm{dim}(\phi)=d<\infty$. (a) For $\alpha=1$, assuming $n\geq 32e^{2}\log(2/\delta)$, the following bound holds with probability at least $1-\delta$:
$\Bigl{|}\,\log\bigl{(}\mathrm{Vendi}_{1}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}\bigr{)}-\log\bigl{(}\mathrm{Vendi}_{1}\bigl{(}P_{X}\bigr{)}\bigr{)}\,\Bigr{|}$
[…]
(b) For every $21<\alpha<$ 2 and $n\geq 2+8\log(1/\delta)$, the following bound holds with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{Vendi}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{Vendi}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}$
[…]
Proof. We defer the proof to the Appendix.
Therefore, assuming a finite feature map $d<\infty$ and given an entropy order $21\leq\alpha<$ 2, the above results indicate the convergence of the Vendi score to the underlying population Vendi given $n=O(d^{2-\alpha})$ samples. Observe that this result is consistent with our numerical observations of the convergence of Vendi score using the finite-dimension cosine similarity kernel in Figure 1.
5 Truncated Vendi Score and its Estimation via Proxy Kernels
Corollaries 1, 2 demonstrate that if the Vendi score order $\alpha$ is greater than $2$ or the kernel feature map dimension $d$ is finite, then the Vendi score can converge to the population Vendi with $n=O(d)$ samples. However, the theoretical results do not apply to an order $21\leq\alpha<$ 2 when the kernel map dimension is infinite, e.g. the original order-1 Vendi score [Friedman and Dieng, 2023] with a Gaussian kernel. Our numerical observations indicate that a standard sample size below 20000 could be insufficient for the convergence of order-1 Vendi score (Figure 1). To address this gap, here we define the truncated Vendi score by truncating the eigenspectrum of the kernel matrix, and then show that the existing kernel approximation algorithms for Vendi score concentrate around this modified Vendi score.
@@ 5 Truncated Vendi Score and its Estimation via Proxy Kernels
Consider the normalized kernel matrix $\frac{1}{n}K$ of samples $x_{1},\ldots,x_{n}$. Then, for an integer parameter $t\geq 1$, consider the top-$t$ eigenvalues of $\frac{1}{n}K$: $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{t}$. Define $S_{t}=\sum_{i=1}^{t}\lambda_{i}$ and consider the truncated probability sequence $[{\lambda}^{\scriptscriptstyle\text{trunc}}_{1},\ldots,{\lambda}^{\scriptscriptstyle\text{trunc}}_{t}]$:
${\lambda}^{\scriptscriptstyle\text{trunc}}_{i}=\lambda_{i}+\frac{1-S_{t}}{t}\quad\text{for $\;i=1,\ldots,t$}$
We define the order-$\alpha$ $t$-truncated Vendi score as
$\mathrm{Vendi}_{\alpha}^{(t)}(x_{1},\ldots,x_{n}):=\exp\Bigl{(}\frac{1}{1-\alpha}\log\Bigl{(}\sum_{i=1}^{t}{\lambda}^{{\scriptscriptstyle\text{trunc}}^{\large\alpha}}_{i}\Bigr{)}\Bigr{)}$
Notably, for order $\alpha=1$, the $t$-truncated Vendi score is:
$\mathrm{Vendi}_{1}^{(t)}(x_{1},\ldots,x_{n}):=\exp\Bigl{(}\sum_{i=1}^{t}{\lambda}^{\scriptscriptstyle\text{trunc}}_{i}\log\frac{1}{{\lambda}^{\scriptscriptstyle\text{trunc}}_{i}}\Bigr{)}$
Remark 1.
The above definition of $t$-truncated Vendi score leads to the definition of $t$-truncated population Vendi $\mathrm{Vendi}_{\alpha}^{(t)}(P_{X})$, where the mentioned truncation process is applied to the eigenspectrum of the population kernel covariance matrix $\widetilde{C}_{X}$. Note that the truncated Vendi score is a statistic and function of random samples $x_{1},\ldots,x_{n}$, whereas the truncated population Vendi is deterministic and a characteristic of the population distribution $P_{X}$.
According to Definition 2, we find the probability model with the minimum $\ell_{2}$-norm difference from the $t$-dimensional vector $[\lambda_{1},\ldots,\lambda_{t}]$ including only the top-$t$ eigenvalues. Then, we use the order-$\alpha$ entropy of the probability model to define the order-$\alpha$ $t$-truncated population Vendi. Our next result shows that this population quantity can be estimated using $n=O(t)$ samples by $t$-truncated Vendi score for every kernel function.
@@
Consider the setting in Theorem 1. Then, for every $n\geq 2+8\log(1/\delta)$, the difference between the $t$-truncated population Vendi and the empirical $t$-truncated Vendi score of samples $x_{1},\ldots,x_{n}$ is bounded with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{Vendi}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{Vendi}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}$
[…]
Proof. We defer the proof to the Appendix. As implied by Theorem 2, the $t$-truncated population Vendi can be estimated using $O(t)$ samples, i.e. the truncation parameter $t$ plays the role of the bounded dimension of a finite-dimension kernel map. Our next theorem shows that the Nyström method [Friedman and Dieng, 2023] and the FKEA method [Ospanov et al., 2024b] for reducing the computational costs of Vendi scores have a bounded difference with the truncated population Vendi.
@@
Consider the setting of Theorem 1. (a) Assume that the kernel function is shift-invariant and the FKEA method with $t$ random Fourier features is used to approximate the Vendi score. Then, for every $\delta$ satisfying $n\geq 2+8\log(1/\delta)$, with probability at least $1-\delta$:
$\Bigl{|}\mathrm{FKEA}\text{-}\mathrm{Vendi}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{Vendi}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\Bigr{|}$
$\,\leq\sqrt{\frac{128\max\{1,t^{2-\alpha}\}\log\bigl{(}3/\delta\bigr{)}}{\min\{n,t\}}}$
@@
$\Bigl{|}\mathrm{Nystrom}\text{-}\mathrm{Vendi}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{Vendi}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\Bigr{|}$
$\ \leq\mathcal{O}\Bigl{(}\sqrt{\frac{\max\{1,t^{2-\alpha}\}\log\bigl{(}2/\delta\bigr{)}t\tau^{2}\log(n)^{2}}{n}}\Bigr{)}$
Proof. We defer the proof to the Appendix.